OpenStudy (anonymous):

Teresa and Julia are among 10 students who have applied for a trip to Washington, D.C. Two students from the group will be selected at random for the trip. What is the probability that Teresa and Julia will be the 2 students selected?

OpenStudy (thurscon):

This is a combination problem. nCr is the general formula n! / ( (n - r)!r!) n = 10 r = 2 10! / [8! 2! ] 10*9/2 = 45 So the probability they will be the students picked is 1/45.

OpenStudy (anonymous):

I get how you got 45, but where did the 1 come from?

OpenStudy (thurscon):

1 just indicates that for every 45 possibilities of student combinations, the combination of those two girls will only occur (in theory) once.

OpenStudy (thurscon):

It's just the reduced version of 2/90. 2/90=1/45

OpenStudy (anonymous):

Why is it 2/90

OpenStudy (kropot72):

This problem is sampling without replacement. The hypergeometric distribution applies.

OpenStudy (thurscon):

There are 10 students to pick from the first time. To determine the second student, there will be 9 left to choose from. So 10*9 = 90. Since there are two girls in the problem, it is 2/90.

OpenStudy (anonymous):

oh i see. Thanks a Bunch :D (Im reviewing for state testing tomorrow)

OpenStudy (thurscon):

Sure, no problem. Best of luck!

OpenStudy (thurscon):

Not quite sure where you got 2*10*9 from. Even using the hg calculation, it would be (2 choose 2) * (8 choose 0) / (10 choose 2) .... which results in 1/45

OpenStudy (kropot72):

The hypergeometric calculation gives a probability of choosing Teresa and Julia as follows:\[P(T+J)=\frac{\left(\begin{matrix}2 \\ 2\end{matrix}\right)\left(\begin{matrix}8 \\ 0\end{matrix}\right)}{\left(\begin{matrix}10 \\ 2\end{matrix}\right)}=\frac{2}{10\times 9}=\frac{1}{45}\] Sorry for mistake in previous posting (now deleted).